# A Physics Preserving Neural Network Based Approach for Constitutive Modeling of Isotropic Fibrous Materials

### v4

@@ 3.2 Numerical Constitutive Model
At this point, we comment on the accuracy of the trained ICNN for individual components of the desired outputs. We use the normalized MSE, computed as $\text{NMSE}(y)=\text{MSE}(y)/\text{mean}(y_{\text{true}}^{2})$, as the measure of deviation from the true value (numerical solution). To start, we present the true vs predicted values of the strain energy density function for the test samples in Figure 4. In this figure, the ICNN-predicted strain energy density corresponding to the numerical solution is shown via scatters. The red line indicates the ideal match between true and predicted value (i.e., slope = 1.) The computed NMSE for strain energy density is approximately 0.15%, indicating exceptional predictive accuracy. Similarly, the true vs predicted values for the unique components of the Cauchy stress tensor are shown in Figure 5. The average NMSE for the stress tensor components is 0.815%, with a maximum of 1.25% from $\sigma_{22}$ component.
@@
Figure 5: True (MuMFiM) vs ICNN-predicted values for unique components of Cauchy stress tensor for various deformation gradients ($F$) in test dataset. The component averaged value of normalized MSE is 0.815%. The result is based on ICNN trained with the $\mathcal{L}^{2}$ loss function.
@@
The mean normalized error increased to 5.4% for the stiffness tensor components shown in Figure 6. The maximum deviation was recorded for $\mathop{\mathbb{D}}_{1312}$ at 23.35%, which is also reflected in the figure as a visible drift from the ideal fit. The loss of accuracy for stiffness is not necessarily surprising. Two potential reasons are: (i) an amplification of noise due to numerical differentiation techniques used to compute the derivative quantities, and (ii) the near-constant gradient of the activation functions at large inputs. The modeling is further complicated by the fact that the underlying data relates to a stochastic fibrous material, which is not guaranteed to have a smooth energy landscape. Despite these challenges, the machine-learned constitutive model is able to faithfully reproduce results on a single RVE (Figure 7) and on an exemplar Facet Capsular Ligament geometry (Figure 8).
@@
Figure 6: True (MuMFiM) vs ICNN-predicted values of the unique components of stiffness tensor for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{2}$ loss function. The average normalized MSE score is 5.4% (maximum: 23%, minimum: 0.03%).
@@ 3.3 Single RVE Results
Figure 7: Comparison of Cauchy stress ($\sigma$) vs deformation gradient ($F$) between a single MuMFiM RVE and ICNN-predicted values for each considered composite loss function undergoing pure shear (left), uniaxial deformation (center), and isochoric deformation (right). For each case, the ICNN neural network with the $\mathcal{L}^{2}$ composite loss function maintains accuracy outside the training regime ($0.85\leq\lambda\leq 1.15$).
@@
We observe that for each case, increasing the number of derivative terms in the composite loss function increases the regime where the ICNN-based model accurately predicts the constitutive response. This is most pronounced in uniaxial extension. We additionally observe that with the $\mathcal{L}^{2}$ composite loss function, the ICNN-based model maintains accuracy outside the training regime ($0.85\leq\lambda\leq 1.15$). These results indicate that it is beneficial to include higher than second derivatives into the training process. However, obtaining numerical or experimental values of these derivatives presents a considerable challenge since derivatives tend to amplify noise. To further emphasize the utility of the ICNN-based models undergoing complex multiaxial load states, we apply them in a 3D finite element simulation of the Facet Capsular Ligament.
@@ 3.4 Case Study: Uniaxial Tension of Facet Capsular Ligament
To demonstrate the efficacy of the ICNN-based model, we present a comparison of the results of a boundary value problem with a 3D model of facet capsular ligament geometry (FCL). To this end, we consider “bone” boundary conditions on the left and right surfaces of the FCL model. In this setup, the left surface is held encastre, and the right surface is fixed in the y and z directions and displaced in the x direction. The meshed geometry and boundary conditions are shown schematically in Figure 8(a). The remaining boundaries are in a traction-free condition. We solve this problem in two ways: First, we use the MuMFiM package to obtain a fully numerical solution. Next, we solve the same problem with trained ICNN as the material model. In other words, we replace the ’microscale module’ in Figure 1a with an ICNN-based framework. We treat the MuMFiM solution as the reference result for this comparison. A more detailed study of this problem is available in [1].
@@
The deformed configurations at approximately 70% global strain obtained from ICNN and MuMFiM-based simulations are shown in Figure 8b and 8c, respectively. The color represents the true stress in the elements along the loading direction. Visually comparing these results, it is immediately clear that the ICNN-based approach closely replicates the predictions of MuMFiM, including the shear-band-like phenomena appearing due to the geometry of the ligament. We also show the result obtained with a calibrated compressible neo-Hookean model in 8d. This model predicts a more diffused band with a far smaller maximum stress. Although neo-Hookean is the simplest material model for hyperelastic behavior, it captures the essence of the difficulties associated with classical continuum models in the context of fibrous network materials. The small parameter space of classical models and strict assumptions regarding the shape of the strain energy density functions fail to capture the complexities arising from the kinematics of the fibers in fibrous materials. A neural network model such as the one developed here thus presents a trade-off between explainability and generality.
@@
The stress-strain response of the facet capsular ligament extracted from the MuMFiM and ICNN-based simulation are presented in Figure 9. Both approaches predict exponential stiffening, which is a hallmark behavior of soft biological tissues [41, 30, 42]. The neo-Hookean model calibrated at the small strain limit is also shown for completeness. The results indicate that the ICNN-based method retains reasonable accuracy up to a very large strain, while reducing the computational resource requirement by orders of magnitude—from GPU-accelerated super-computers to consumer-grade laptops. The inset of Figure 9 shows the measured relative error of the ICNN result for stress with respect to MuMFiM. The maximum relative error is in the small-strain regime and upward of 50%. The deviation, however, falls rapidly below 20% at approximately 10% strain. The discrepancy in the small strain regime is related to a similar deviation in the predicted stiffness, where the true values are close to zero (Figure 6) and the relative dominance of numerical noise when the true value is small. In fibrous network materials such as biological tissues, however, the large strain regime is the principal concern [9]. Given the computational cost for large-scale simulation of fibrous materials and the technical difficulties required to develop efficient algorithms such as those employed in [1], the current model presents an excellent pathway for obtaining approximate results involving the deformation of fibrous materials. Although our work does not obviate the need for large-scale numerical simulations, it does provide a pathway for those in the community to make use of results from those expensive simulations with modest computational resources.
@@
Figure 8: (a) Geometry and boundary conditions for numerical simulation of facet capsule ligament (FCL) under uniaxial tension along the x-axis. (b) Deformed configuration at 70% strain based on large scale MuMFiM/FE2 simulation. (c) Deformed configuration at same strain using standard finite element method with trained ICNN ($\mathcal{L}^{2}$ variant) as the constitutive descriptor, and (d) with calibrated neo-Hookean material model. The color indicates the true stress in the elements along the loading direction.
@@ 3.5 Limitations
    Limited availability of training data: Our model relies on the higher-order derivatives of the energy function to achieve superior predictive accuracy for quantities such as stiffness tensor components. These datasets are generally not readily available, especially for novel materials. Under these limiting circumstances, we suggest making use of alternative models that rely on available datasets only while being mindful of the loss of accuracy for higher-order quantities.
@@
    Limited accuracy in the small strain regime: Our model does not enforce $\mathcal{W}=0$ for $\mathbf{F}=\mathbf{I}$ due to the model architecture, specifically, the same constraints that ensure the model convexity. However, the optimizer can adjust the bias in pass-through layers without any constraints to push $\mathcal{W}$ arbitrarily close to zero. The trained model indeed shows such convergence for strain energy density and stress (Figure 4 and 5). Additionally, the numerical derivatives used to compute the higher-order quantities amplify the noise. We speculate these two factors combined produce this discrepancy.
@@
We hope the published code and dataset will increase interest in this approach and lead to more general models encompassing other microstructural aspects such as fiber prealignment and flocculation.
@@ 4 Conclusion
We show that constitutive constraints alone are not enough to construct accurate models of stress and stiffness. The contributions of this work stem from: First, the incorporation of the Sobolev training protocol, which involves including higher-order derivatives in the loss function, significantly improves the model’s prediction accuracy for stress and stiffness. Second, the trained model, when used as a replacement for the subscale segment of multiscale procedure such as FE2 or as a material model in standard finite element procedures, provides a quantitatively similar solution to practical problems of interest at a fraction of the computational cost. Despite their accuracy, performing FE2 multiscale simulations requires significant computational resources that are out of reach for most biomechanical researchers. Our material model provides a bridge between the computational demands of multiscale simulations and the typical accuracy needed to derive important biomechanical insights. We exemplify this through the analysis of a Facet Capsular Ligament and believe that others can benefit from applying this method to their research.

### v5

@@ 3.2 Numerical Constitutive Model
At this point, we comment on the accuracy of the trained ICNN for individual components of the desired outputs. We use the normalized MSE, computed as $\text{NMSE}(y)=\text{MSE}(y)/\text{mean}(y_{\text{true}}^{2})$, as the measure of deviation from the true value (numerical solution).
@@
Here, we present the true vs predicted values of the strain energy density for the test samples in Figure 4 for various choices of the loss function and input invariants. In this figure, the ICNN-predicted strain energy density corresponding to the numerical solution is shown via scatters. The red line indicates the ideal match between true and predicted value (i.e., slope = 1.) The computed NMSE for strain energy density is approximately 0.15%, indicating exceptional predictive accuracy. Further, the accuracy of the strain energy is similar for all choices of loss function, and input invariants.
The corresponding results for the unique components of Cauchy stress tensors are presented in Figure 5 which reaffirms the conclusion drawn from Figure 3, viz, the predictive accuracy for stresses is similar for $\mathcal{L}^{1}$ and $\mathcal{L}^{2}$ when the input array consists of all isotropic invariants. The average NMSE for the stress tensor components is 0.815%, with a maximum of 1.25% from $\sigma_{22}$ component with $\mathcal{L}^{2}(I_{1},I_{2},I_{3})$.
However, unlike energy, omitting $I_{2}$ as an input degrades the predictive accuracy for stress for the material under consideration. This is seen as a wide band in Figure 5c. Omitting $I_{3}$ as an input does not impact the predictive accuracy of the stress (Figure 5d). This indicates that the material can be considered nearly incompressible. Detailed error measures for the stress components are given in SI (Fig 3).
@@
Figure 5: True (MuMFiM) vs ICNN-predicted values for unique components of Cauchy stress tensor for various deformation gradients ($F$) in the test dataset. The variant of the loss function used and invariants of $\mathbf{C}$ included in the input array are indicated in sub-figure titles.
We observe a similar trend for the components of stiffness tensor visualized in Figure 6-10. For the given ICNN architecture, the prediction accuracy is best of $\mathcal{L}^{2}$ variant with at least $I_{1}$ and $I_{2}$ included in the input array. Here, the mean normalized error is 5.4% for the stiffness tensor components. The maximum deviation is for $\mathop{\mathbb{D}}_{1312}$ at 23.35%, which is also reflected as a visible drift from the ideal fit in Figure 10 and 9.
@@
This loss of accuracy for stiffness is not necessarily surprising. Two potential reasons are (i) poor signal-to-noise ratio at the small strain limit in the dataset due to higher order derivatives evaluated using numerical methods and (ii) the near-constant gradient of the activation functions at large inputs. The modeling is further complicated by the fact that the underlying data relates to a stochastic fibrous material, which is not guaranteed to have a smooth energy landscape. Despite these challenges, the machine-learned constitutive model is able to faithfully reproduce results on a single RVE (Figure 11) and on an exemplar FCL geometry (Figure 12).
Other variants of the model, e.g., $\mathcal{L}^{0},\mathcal{L}^{1}$ and $\mathcal{L}^{2}(I_{1}\&I_{3})$ generally performed worse for material under consideration as shown in Figure 6, 7, and 8 respectively. The detailed error measures for the variants are available in SI (Fig 4).
@@
Figure 6: True (MuMFiM) vs ICNN-predicted values of the unique components of stiffness tensor for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{0}$ loss function. The input array includes all isotropic invariants of $\mathbf{C}$.
Figure 7: True (MuMFiM) vs ICNN-predicted values of the unique components of stiffness tensor for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{1}$ loss function. The input array includes all isotropic invariants of $\mathbf{C}$.
Figure 8: True (MuMFiM) vs ICNN-predicted values of the unique components of stiffness tensor for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{2}$ loss function with $I_{1}$ and $I_{3}$ as input.
Figure 9: True (MuMFiM) vs ICNN-predicted values of the unique components of stiffness tensor for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{2}$ loss function with $I_{1}$ and $I_{2}$ as input.
Figure 10: True (MuMFiM) vs ICNN-predicted values of the unique components of stiffness tensor for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{2}$ loss function. The input array includes all isotropic invariants of $\mathbf{C}$.
@@ 3.3 Single RVE Results
Figure 11: Comparison of Cauchy stress ($\sigma$) vs deformation gradient ($F$) between a single MuMFiM RVE and ICNN-predicted values for each considered composite loss function undergoing pure shear (left), uniaxial deformation (center), and isochoric deformation (right). For each case, the ICNN neural network with the $\mathcal{L}^{2}$ composite loss function maintains accuracy outside the training regime ($0.85\leq\lambda\leq 1.15$). The ICNN was trained with all isotropic invariants of $\mathbf{C}$
@@
We observe that for each case, increasing the number of derivative terms in the composite loss function increases the regime where the ICNN-based model accurately predicts the constitutive response. This is most pronounced in uniaxial extension. We additionally observe that with the $\mathcal{L}^{2}$ composite loss function, the ICNN-based model maintains accuracy outside the training regime ($0.85\leq\lambda\leq 1.15$). These results indicate that it is beneficial to include higher than second derivatives into the training process. To further emphasize the utility of the ICNN-based models undergoing complex multiaxial load states, we apply them in a 3D finite element simulation of the FCL.
@@ 3.4 Case Study: Uniaxial Tension of Facet Capsular Ligament
To demonstrate the efficacy of the ICNN-based model, we present a comparison of the results of a boundary value problem with a 3D model of a FCL. The shape of the FCL cross-section was obtained from [49] and is representative of a human Cervical FCL. The cross-section is uniformly extruded into the plane with a thickness of 0.5mm using Siemens NX CAD. The bounding box around the ligament is 11.84mm by 8.56mm in the x and z directions. This model was meshed and model attributes were assigned in Simmetrix SimModeler [50]. The mesh has 63,821 linear tetrahedral elements. Meshes were converted to the PUMI native format and model attributes were exported as a model-traits YAML file which are read by MuMFiM [51, 52].
For the case presented herein, Delaunay networks are constructed by the procedure in 2.1.4 that serve as RVE for MuMFiM. Fiber orientations are uniformly distributed in the initial configuration (no preferential alignment). They have average density defined as the total fiber length per unit volume of $186.12$ and a mean fiber length of $0.216$.
@@
We consider “bone” boundary conditions on the left and right surfaces of the FCL model. In this setup, the left surface is held encastre, and the right surface is fixed in the $y$ and $z$ directions and displaced in the $x$ direction. The meshed geometry and boundary conditions are shown schematically in Figure 12(a). The remaining boundaries are in a traction-free condition. We solve this problem in two ways: First, we use the MuMFiM package to obtain a fully numerical solution. Next, we solve the same problem with trained ICNN as the material model. In other words, we replace the ’microscale module’ in Figure 1a with an ICNN-based framework. We treat the MuMFiM solution as the reference result for this comparison. A more detailed study of this problem is available in [1].
@@
The deformed configurations at approximately 70% global strain obtained from ICNN and MuMFiM-based simulations are shown in Figure 12b and 12c, respectively. The color represents the true stress in the elements along the loading direction. Visually comparing these results, it is immediately clear that the ICNN-based approach closely replicates the predictions of MuMFiM, including the shear-band-like phenomena appearing due to the geometry of the ligament. We also show the result obtained from a compressible neo-Hookean model that has been calibrated on the small strain data from MuMFiM in 12d. This model predicts a more diffused band with a far smaller maximum stress. Although neo-Hookean is the simplest material model for hyperelastic behavior, it captures the essence of the difficulties associated with classical continuum models in the context of fibrous network materials. The small parameter space of classical models and strict assumptions regarding the shape of the strain energy density functions fail to capture the complexities arising from the kinematics of the fibers in fibrous materials. A neural network model such as the one developed here thus presents a trade-off between explainability and generality.
@@
The stress-strain response of the FCL extracted from the MuMFiM and ICNN-based simulation are presented in Figure 13. Both approaches predict exponential stiffening, which is a hallmark behavior of soft biological tissues [53, 37, 54]. The neo-Hookean model calibrated with MuMFiM data at a small strain limit does not show similar exponential stiffening. The results indicate that the ICNN-based method retains reasonable accuracy up to a very large strain while reducing the computational resource requirement by orders of magnitude—from GPU-accelerated super-computers to consumer-grade laptops.
The inset of Figure 13 shows the measured relative error of the ICNN result for stress with respect to MuMFiM. The maximum relative error is in the small-strain regime and upward of 50%. The deviation, however, falls rapidly below 20% at approximately 10% strain. The discrepancy in the small strain regime is likely related to a similar deviation in the predicted stiffness at the small strain limit (Figure 10). We suspect the application of numerical noise at this limit is behind this discrepancy. In fibrous network materials such as biological tissues, however, the large strain regime is the principal concern [9]. Given the computational cost for large-scale simulation of fibrous materials and the technical difficulties required to develop efficient algorithms such as those employed in [1], the current model presents an excellent pathway for obtaining approximate results involving the deformation of fibrous materials. To put the cost savings into perspective, the multiscale analysis requires one RVE solve per element in each iteration of the macroscale solver. In this case, we used a line-search method that does three evaluations to compute a valid search distance with an average of six iterations per load step. For the 60 load steps in the FCL model, MuMFiM must perform nearly 69 million evaluations of the discrete fiber network RVEs (63821 elements * 6 line search per step * 3 iterations per line search * 60 steps). Unsurprisingly, these RVE evaluations account for the majority of the MuMFiM solve time. The dataset that was used to train the model in this paper only required 91,961 evaluations of the discrete RVE. Although our work does not obviate the need for large-scale numerical simulations, it does provide a pathway for those in the community to make use of results from those expensive simulations with modest computational resources.
@@
Figure 12: (a) Geometry and boundary conditions for numerical simulation of FCL under uniaxial tension along the x-axis. This model has 63,821 linear tetrahedral elements. (b) Deformed configuration at 70% strain based on large scale MuMFiM/FE2 simulation. (c) Deformed configuration at same strain using standard finite element method with trained ICNN ($\mathcal{L}^{2}$ variant) as the constitutive descriptor, and (d) with calibrated neo-Hookean material model. The color indicates the true stress in the elements along the loading direction normalized by the small strain stiffness of the material, $E_{o}$. A different color bar is used for the Neo-Hookean model to emphasize that both the magnitude and structure of the field are different for the Neo-Hookean material. The reader is referred to Figure 9 for additional insight into the magnitude of the difference.
@@ 3.5 Limitations
    Limited availability of training data: Our model relies on the higher-order derivatives of the energy function to achieve superior predictive accuracy for quantities such as stiffness tensor components. These datasets are generally not readily available, especially for novel materials. Under these limiting circumstances, we suggest employing $\mathcal{L}^{0}$ or $\mathcal{L}^{1}$ loss function variant or models described in [25, 17] while being mindful of the loss of accuracy for higher-order quantities.
@@
    Limited accuracy in the small strain regime: Our model does not enforce $\mathcal{W}=0$ for $\mathbf{F}=\mathbf{I}$ due to the model architecture, specifically, the same constraints that ensure the model convexity. However, the optimizer can adjust the bias in pass-through layers without any constraints to push $\mathcal{W}$ arbitrarily close to zero. The trained model indeed shows such convergence for strain energy density and stress (Figure 4 and 5). We also observe relatively poor predictive accuracy of stiffness tensor components at the small strain limit, which may be a cause of concern for specific use cases.
@@
We hope the published code and dataset will increase interest in this approach and lead to more general models encompassing other microstructural aspects, such as fiber alignment and flocculation.
@@ 4 Conclusion
We show that constitutive constraints alone are not enough to construct accurate models of stress and stiffness. The contributions of this work stem from: First, the incorporation of the Sobolev training protocol, which involves including higher-order derivatives in the loss function, significantly improves the model’s prediction accuracy for stress and stiffness. Second, the trained model, when used as a replacement for the subscale segment of multiscale procedures such as FE2 or as a material model in standard finite element procedures, provides a quantitatively similar solution to practical problems of interest at a fraction of the computational cost. Despite their accuracy, performing FE2 multiscale simulations requires significant computational resources that are out of reach for most biomechanical researchers. Our material model provides a bridge between the computational demands of multiscale simulations and the typical accuracy needed to derive important biomechanical insights. We exemplify this through the analysis of a FCL and believe that others can benefit from applying this model and method to their research.